OpenStudy (anonymous):

y+12=5y-4? I got y=-10 just checking

OpenStudy (adamaero):

nope

OpenStudy (anonymous):

its 4

OpenStudy (anonymous):

why ? D:

OpenStudy (anonymous):

i can show you might work

OpenStudy (anonymous):

my

OpenStudy (mitu12):

4

OpenStudy (anonymous):

k

OpenStudy (anonymous):

how u guys got 4 ?

OpenStudy (anonymous):

y+12=5y-4 +4 +4 y+16=5y -y -y 16=4y then divide

OpenStudy (mitu12):